OpenStudy (anonymous):

evaluate cot ( Cos^-1( -15/17)) enter your answer as a fraction using the slash bar (/)

OpenStudy (anonymous):

inverse cosine means it's going to give you the angle that yields that ratio of sides. Cos is adjacent over hypotenuse, so draw a triangle with -15 as the adjacent side and 17 as the hypotenuse: 

OpenStudy (anonymous):

Now, simply use the pythagorean theorem to solve for the third side.

OpenStudy (anonymous):

And then use the definition of cot = adjacent over opposite.

OpenStudy (anonymous):

i tried but gave me 6/4 I'm lost :/

OpenStudy (anonymous):

Show me the math you did for solving for the third side.

OpenStudy (anonymous):

Pythagorean theorem: a^2 + b^2 = c^2

OpenStudy (anonymous):

is in the trash :/

OpenStudy (paxpolaris):

 Finding the third side: \[\Large x=\sqrt{17^2-15^2}\]

OpenStudy (anonymous):

lols, okay. Well try it again.

OpenStudy (anonymous):

i don't have the calculator anymore :( and im helping my bro with his Work

OpenStudy (anonymous):

Use google as a calculator =) Or use wolframalpha.com

OpenStudy (anonymous):

hint: to do square roots, type sqrt(whatever)

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

Don't pout at me. Gimme some effort.

OpenStudy (paxpolaris):

\[\Large 17^2 =\left( 15+2 \right)^2=225+60+4\]

OpenStudy (anonymous):

289 but i need it in fractions

OpenStudy (paxpolaris):

\[x= \sqrt{289-225}\]

OpenStudy (anonymous):

-_-

OpenStudy (paxpolaris):

x is the third side.... first find it the you can solve for cotangent

OpenStudy (anonymous):

still dont get it -_______________________-